OpenStudy (anonymous):

What is the exact value of the expression 3the square root of 11. + the square root of 6. - 9the square root of 11.? -5the square root of 17. -6the square root of 11. + the square root of 6. -3the square root of 11. + the square root of 6. -5the square root of 6.

OpenStudy (anonymous):

is it\[3\sqrt{11} + \sqrt{6} - 9\sqrt{11}?\]

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

@Crixpack

OpenStudy (anonymous):

Tell me what you think about it so I can help you out

OpenStudy (anonymous):

I'm confused about it all on this one

OpenStudy (anonymous):

I'm gonna help you out a but and rearrange them like this: \[3\sqrt{11} - 9\sqrt{11} +\sqrt{6}\] Do you see any way this can be simplified?

OpenStudy (anonymous):

Not really

OpenStudy (anonymous):

\[\sqrt{11}\] is just a number, think of it like x, so the first two terms are basically 3x - 9x, so what does that give us

OpenStudy (anonymous):

-6 @Crixpack

OpenStudy (anonymous):

not just -6, -6 what

OpenStudy (anonymous):

-6x

OpenStudy (anonymous):

now what?

OpenStudy (anonymous):

right, now for our equation its not x, but \[\sqrt{11}\]

OpenStudy (anonymous):

so what do we have

OpenStudy (anonymous):

do you want me to find the square root of 11?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

what do you want me to do now?

OpenStudy (anonymous):

I just want you to simplify the equation and keep the roots as they are

OpenStudy (anonymous):

\[(3\sqrt{11} - 9\sqrt{11}) +\sqrt{6}\] just simplify the part in the brackets, and keep it in terms of \[\sqrt{11}\]

OpenStudy (anonymous):

so is the answer B?

OpenStudy (anonymous):

Indeed it is

OpenStudy (anonymous):

Awesome! Thanks for helping me figure it out!

OpenStudy (anonymous):

No problem :)